# Data Efficient Behavior Cloning for Fine Manipulation via Continuity-based Corrective Labels

### v1

@@ Abstract
We consider imitation learning with access only to expert demonstrations, whose real-world application is often limited by covariate shift due to compounding errors during execution. We investigate the effectiveness of the Continuity-based Corrective Labels for Imitation Learning (CCIL) framework in mitigating this issue for real-world fine manipulation tasks. CCIL generates corrective labels by learning a locally continuous dynamics model from demonstrations to guide the agent back toward expert states. Through extensive experiments on peg insertion and fine grasping, we provide the first empirical validation that CCIL can significantly improve imitation learning performance despite discontinuities present in contact-rich manipulation. We find that: (1) real-world manipulation exhibits sufficient local smoothness to apply CCIL, (2) generated corrective labels are most beneficial in low-data regimes, and (3) label filtering based on estimated dynamics model error enables performance gains. To effectively apply CCIL to robotic domains, we offer a practical instantiation of the framework and insights into design choices and hyperparameter selection. Our work demonstrates CCIL’s practicality for alleviating compounding errors in imitation learning on physical robots.
@@ I Introduction
We focus on the domain of fine manipulation, an area that encompasses a wide range of real-world applications, from manufacturing to assistive robotics [19, 20, 21]. Fine manipulation tasks are notably challenging even for human experts due to their demand for precise control and susceptibility to failure from minor inaccuracies, resulting in limited data availability. However, learning from demonstration is often preferred over model-based methods for many fine manipulation tasks with hard-to-define cost functions (e.g., stitching wounds). Thus, the inherent difficulties of fine manipulation underscore a critical need for strategies to mitigate compounding errors in imitation learning, making it an ideal testing ground.
@@ IV-B CCIL’s Assumptions on Local Lipschitz Continuity
Figure 4: (a) The 20% dataset for the GraspCube task reveals two distinct clusters of corrective labels. (b) The green cluster mainly corresponds to labels where the cube is being manipulated, and the blue cluster to arm-free space motion. (c) Policies trained using just the blue cluster (low label rejection threshold) are more successful compared to those trained with the green cluster or both (high label rejection threshold.
@@ VII Acknowledgement
This work was (partially) funded by the National Science Foundation NRI (#2132848) and CHS (#2007011), DARPA RACER (#HR0011-21-C-0171), the Office of Naval Research (#N00014-17-1-2617-P00004 and #2022-016-01 UW), Amazon and NSF Robust Intelligence Grant (#2212310).

### v3

@@ Abstract
We consider imitation learning with access only to expert demonstrations, whose real-world application is often limited by covariate shift due to compounding errors during execution. We investigate the effectiveness of the Continuity-based Corrective Labels for Imitation Learning (CCIL) framework in mitigating this issue for real-world fine manipulation tasks. CCIL generates corrective labels by learning a locally continuous dynamics model from demonstrations to guide the agent back toward expert states. Through extensive experiments on insertion and fine grasping tasks, we provide the first empirical validation that CCIL can significantly improve imitation learning performance despite discontinuities present in contact-rich manipulation. We find that: (1) real-world manipulation exhibits sufficient local smoothness to apply CCIL, (2) generated corrective labels are most beneficial in low-data regimes, and (3) label filtering based on estimated dynamics model error enables performance gains. To effectively apply CCIL to robotic domains, we offer a practical instantiation of the framework and insights into design choices and hyperparameter selection. Our work demonstrates CCIL’s practicality for alleviating compounding errors in imitation learning on physical robots.
See more: https://personalrobotics.github.io/CCIL/
@@ I Introduction
We focus on the domain of fine manipulation, an area that encompasses a wide range of real-world applications, from manufacturing to assistive robotics [19, 20, 21]. Fine manipulation tasks are notably challenging even for human experts due to their demand for precise control and susceptibility to failure from minor inaccuracies, resulting in limited data availability. However, learning from demonstration is often preferred over model-based methods for many fine manipulation tasks with hard-to-define cost functions (e.g., stitching wounds), or hard to model dynamics. Thus, the inherent difficulties of fine manipulation underscore a critical need for strategies to mitigate compounding errors in imitation learning, making it an ideal testing ground.
@@ IV-B CCIL’s Assumptions on Local Lipschitz Continuity
Figure 4: (a) The 20% dataset for the GraspCube task reveals two distinct clusters of corrective labels. (b) The green cluster mainly corresponds to labels where the cube is being manipulated, and the blue cluster to arm-free space motion. (c) Policies trained using just the blue cluster (low label rejection threshold) are more successful compared to those trained with the green cluster or both (high label rejection threshold).
@@ VII Acknowledgement
This work was (partially) funded by the National Science Foundation NRI (#2132848) and CHS (#2007011), DARPA RACER (#HR0011-21-C-0171), the Office of Naval Research (#N00014-17-1-2617-P00004 and #2022-016-01 UW), Amazon and funding from the Toyota Research Institute through the University Research Program 3.0.